OpenStudy (akgorgeous):

ORIGINAL PROBLEM- ( A scientist is studying the growth of a particular species of plant. He writes the following equation to show the height of the plant f(n), in cm, after n days: f(n) = 8(1.05)n ) MY QUESTION- what is the y intercept of the equation ( f(n) = 8(1.05)n )?

OpenStudy (akgorgeous):

@retirEEd

OpenStudy (johnsmith18):

The y-intercept is 0 since it is the plant's initial height.

OpenStudy (retireed):

Actual the equation as written is misleading.... f(n) = 8(1.05)n should be written as f(n) = 8(1.05)^n I believe the n is an exponent When n is zero the 1.05^0is equal to 1 thus f(n) = 8 times 1 equals 7